# Unveiling Language-Specific Features in Large Language Models via Sparse Autoencoders

### v1

@@ Abstract
The mechanisms behind multilingual capabilities in Large Language Models (LLMs) have been examined using neuron-based or internal-activation-based methods. However, these methods often face challenges such as superposition and layer-wise activation variance, which limit their reliability. Sparse Autoencoders (SAEs) offer a more nuanced analysis by decomposing the activations of LLMs into sparse linear combination of SAE features. We introduce a novel metric to assess the monolinguality of features obtained from SAEs, discovering that some features are strongly related to specific languages. Additionally, we show that ablating these SAE features only significantly reduces abilities in one language of LLMs, leaving others almost unaffected. Interestingly, we find some languages have multiple synergistic SAE features, and ablating them together yields greater improvement than ablating individually. Moreover, we leverage these SAE-derived language-specific features to enhance steering vectors, achieving control over the language generated by LLMs.
@@
Boyi Deng1thanks: Work done during internships at Alibaba Group., Yu Wan1thanks: Corresponding author, Yidan Zhang1, Baosong Yang1, Fuli Feng2 1Tongyi Lab, Alibaba Group Inc, 2Institute of Dataspace, Hefei, Anhui, China dengboyi@mail.ustc.edu.cn {wanyu.wy,nianjun.zyd}@alibaba-inc.com
@@ 1 Introduction
Large Language Models (LLMs) (OpenAI et al., 2024; Grattafiori et al., 2024; Qwen et al., 2025) exhibit impressive abilities in various domains such as text generation (OpenAI et al., 2024; Grattafiori et al., 2024), instruction following (Zhang et al., 2024a; Lou et al., 2024), and reasoning (Huang and Chang, 2023). Recently, considerable efforts have been made to enhance the multilingual capabilities of LLMs to meet the growing demand for their deployment in multilingual environments (Qin et al., 2024; Huang et al., 2025). For instance, Gemini 1.5 incorporates a variety of multilingual data in its training process and emphasizes its multilingual capabilities (Team et al., 2024a)Yang et al. (2024) claim that Qwen2 supports over 30 languages and achieve great performance on multilingual benchmarks. Moreover, multilingual training data comprises approximately 3% of the training data for Llama 3, and there are also high-quality multilingual instruction-tuning data for 8 languages (Grattafiori et al., 2024). As the significance of multilingual capabilities in LLMs continues to grow, it is crucial to delve into the mechanisms of these capabilities to enhance them further.
@@
Given the advantages of SAEs, we use them to analyze multilingual capabilities in LLMs. Concretely, we start with a preliminary experiment in which we find high activation of some features in certain language, inspired by this, we propose a metric to measure the monolinguality of a feature based on the activation difference across different languages. The results show that some features possess strong monolingual characteristics. Moreover, we believe that these language-specific features are not only related to language-specific tokens, so we experiment on a “code-switching” (Kuwanto et al., 2024; Winata et al., 2023) dataset and find that language-specific features are also closely associated with the language-specific linguistic context. Furthermore, we use directional ablation (Arditi et al., 2024; Ferrando et al., 2024) to “zero out” language-specific features during the forward pass of LLMs, resulting in a loss of capabilities in only certain language. Interestingly, we observe that some language may exhibit more than one specific feature. And these features have a synergistic relationship, meaning ablating these features together results in a significant improvement compared to ablating them individually.
The language-specific features we find are of great monolinguality, so we further leverage them to improve steering vectors (Turner et al., 2024). Concretely, we use language-specific features as gating signals to control steering vectors, and achieve better control over the language generated by LLMs, which valid the practical potential of these language-specific features.
@@
  • We find that ablating language-specific features only significantly decrease the language-specific capabilities of LLMs.
@@ 2 Preliminary
Figure 1: The values of $\nu$, as referenced in Eq. 4 where a larger $\nu$ indicates stronger monolingualism, are reported for the top-4 features and a random feature across various languages in layer 20 of Gemma 2 2B. The values of $\nu$ for the top-4 features are greater than that of a random feature. In most languages, the top-1 feature possesses a significantly larger $\nu$. Additional results for other layers and LLMs are in Appendix C, exhibiting similar patterns.
@@
Flores-200 (Costa-jussà et al., 2022; Goyal et al., 2022) is a parallel corpus that contains translations of English sentences into 200 different languages. We extract a subset called Flores-10, which includes 10 languages [^2] English (en), Spanish (es), French (fr), Japanese (ja), Korean (ko), Portuguese (pt), Thai (th), Vietnamese (vi), Chinese (zh), and Arabic (ar)..
Figure 2: The mean activation of feature 13788 across different languages in layer 10 of Gemma 2 2B. The high mean activation in Chinese suggest that feature 13788 might be related to Chinese.
@@ 3.1 Finding Language-Specific Features
To find language-specific features, we conduct a preliminary experiment by prompting Flores-10 into the LLMs and analyzing the residual stream using SAEs. We find that the mean activation of some features is particularly high for a certain language, while remaining very low for other languages, as illustrated by the example in Figure 2.
Inspired by this, we propose a metric to measure the monolinguality of a feature. Specifically, given a set $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{K}\}$, which contains the residual stream set for a certain layer with $K$ different languages, we calculate the mean activation difference of feature $s$ for a specific language $L$ compared to the other languages as follows:
@@
where $\mathbf{f}_{s}(\mathbf{x})$ is the activation of feature $s$. We calculate $\nu$ for all languages and features and rank them from high to low for each language. The top-ranked features are considered as language specific features.
@@ 3.2 Monolinguality Analysis
We use first 100 data in Flores-10 to calculate $\nu$ for each language. The results are shown in Figure 1. From this figure, we make the following observations. (1) The mean activation of the top-4 features is significantly higher than that of a random feature, which remains close to zero. (2) For most languages, the mean activation of the top features decreases rapidly among the first few, and the mean activation of rank #1 feature is considerably higher than the others. (3) In some languages, the rank #2 feature also shows a substantially large mean activation compared to other features. These results suggest that top-ranked features possess strong monolingual characteristics, and in most scenarios, the top-1 feature suffices in capturing these characteristics. [^3] English is the primary language for most LLMs, and it often exhibits different characteristics compared to other languages (Qin et al., 2024), so we only focus on non-English results in the subsequent sections.
@@
Figure 4: The mean activation values for the French and Korean feature with various noun and prefix combinations. Introducing a different language prefix decreases the original language feature activation of nouns.
Figure 5: The changes in CE loss on texts in the target language and texts in other languages after ablating language-specific features. Ablating language-specific features has a much larger impact on the CE loss of texts in target language compared to texts in other languages. We provide results for Gemma 2 2B here, additional results can be found in Appendix D.
@@ 4 Language-Specific Features Extend Beyond Language-Specific Tokens
In earlier sections, we only evaluates language-specific features on monolingual texts. This raises a natural question: are these language-specific features solely related to language-specific tokens? To explore this, we focus on a phenomenon called “code-switching.” [^4] Code-switching refers to the practice of alternating between two or more languages within a single text (Kuwanto et al., 2024; Winata et al., 2023). Our findings indicate that language-specific features are also related to language-specific linguistic context.
@@ 4.2 Results
We analyze the mean activation values of the Spanish features for Spanish, French and Korean nouns, comparing scenarios with and without Spanish prefixes, as illustrated in Figure 3. Our observations are as follows: (1) Introduction of a Spanish prefix to a French or Korean noun results in higher Spanish feature activation values compared to when the French or Korean nouns stand alone. However, the value is still lower than that of the combination of Spanish prefix and Spanish nouns. (2) The activation value for Spanish features of stand-alone French and Korean nouns remains relatively low across all layers. (3) Both French and Korean nouns with a Spanish prefix show greater increases in Spanish feature activations at deeper layers than at shallower ones. (4) Adding a Spanish prefix results in a larger increase in the Spanish feature for French nouns compared to Korean nouns. These findings suggest that adding a Spanish prefix enhances the Spanish feature activation for non-Spanish nouns, enabling the LLM to process them as if they were “Spanish tokens.” Consequently, this allows the LLM to use these non-Spanish tokens within a consistent language context.
@@
We also analyze the mean activation values of the French and Korean features for corresponding nouns, comparing scenarios with and without Spanish prefixes, as presented in the provided Figure 4. Our observations are as follows: (1) For French and Korean nouns, the original language feature activation is significantly higher when the nouns are standalone than when preceded by a Spanish prefix. (2) Both French and Korean nouns show greater decreases in their original language feature activations at deeper layers than at shallower ones. (3) Adding a Spanish prefix results in a larger decrease in the corresponding feature for French nouns compared to Korean nouns. These findings reveal that introducing a different language prefix, such as Spanish, decreases the original language feature activation of nouns, making them less like nouns from their original language.
@@
The results in Figures 3 and 4 suggest that language-specific features are not solely tied to specific language tokens but are also closely associated with the language-specific linguistic context. This suggests that the linguistic characteristics recognized by the model extend beyond individual words to encompass the contextual environment in which these words appear. Notably, the influence of a Spanish prefix is more pronounced on French nouns than on Korean nouns, potentially due to the linguistic similarities between Spanish and French as Romance languages. This highlights the model’s ability to dynamically adapt its feature activations based on the surrounding linguistic context, effectively reinterpreting non-Spanish tokens within a Spanish framework while diminishing their original language attributes.
@@ 5 Ablating Language-Specific Features Leads to Language-Specific Changes
Figure 6: The change in CE loss for three languages after ablating French Features. Simultaneously ablating multiple French features exhibits a synergistic effect in French, while showing no synergistic effect on other languages. We provide results for Gemma 2 2B here, additional results can be found in Appendix F.
@@ 5.2 Ablation of Language-Specific Features
For each target language, layer, and LLM, we intervene in the inference process of LLMs using Eq. 4 on the top-2 language-specific features of the target language. We then measure the changes in CE loss for both texts in target language and texts in other languages after ablating language-specific features. The results are shown in Figure 5. We observe that: (1) Ablating language-specific features has a much larger impact on the CE loss of texts in target language compared to texts in other languages. (2) For different layers, the changes in CE loss of target language texts vary significantly. These findings suggest that language-specific features play a crucial role in controlling the generation process for target language. Ablating these features from the generation process of LLMs can lead to a loss of only specific language capabilities.
@@ 5.3 Synergistic Language Features
We compare the CE loss for French, Spanish and Japanese when using different numbers of French features for directional ablation. The results are shown in Figure 6. From this, we make the following observations: (1) In some layers, simultaneously ablating the top 2 French features for French significantly impacts the CE loss more than ablating these features individually. (2) In all layers, simultaneously ablating the top 2 French features for Spanish and Japanese results in a CE loss impact approximately equal to the sum of the effects when these features are ablated individually. (3) The changes in CE loss for French are larger than those for Spanish and Japanese. The changes for Spanish are large in some layers, while for Japanese, they are nearly zero across all layers.
@@ 6 Enhancing Steering Vectors Using Language-Specific Features
Having studied the basic characteristic of language-specific features, we now explore how to leverage these features in practice, concretely, we use language-specific features as signals to guide steering vector (Turner et al., 2024; Rimsky et al., 2024; Mayne et al., 2024), in order to control the language in the model.
@@ 6.2 Methods
Figure 8: The y-axis shows the CE loss of LLMs identifying text in original language as if it were in the target language, with lower values indicating better performance. The x-axis shows the impact on non-original language texts, with lower values indicating less impact. “SV” is the “steering vector” method, while “SAE” is our enhanced method. The suffix $k$L indicates results from modifying $k$ consecutive layers. Our method (green) provides a better balance between the metrics.
@@
Table 1: The results of Cross-Lingual Continuation task. Our SAE method can surpass the SV method across both metrics in most cases, achieve a much better balance between the metrics.
@@ 6.3 Results
In Figure 8, we present the results for Adversarial Language Identification. The experiments covers the steering vector both with and without SAE across one, two, and three consecutive layers, denoted as ${1L}$, ${2L}$, and ${3L}$. From the figure, we make the following observations: (1) As the number of modified layers increases, our SAE method achieves better performance in Adversarial Language Identification, whereas the performance of the SV method decreases. This indicates that SAE is more effective at Adversarial Language Identification and more robust when applied to multiple layers. (2) As the number of modified layers increases, the CE loss on other languages also increases. However, the rate of increase with our SAE method is much smaller than with the SV method. In most cases, the CE loss on other languages for SAE applied to three consecutive layers is even lower than that for SV applied to a single layer. These results suggests that our SAE method achieve a better balance between two metrics on Adversarial Language Identification.
@@ 7 Related Works
Multilingual mechanism of LLMs are mainly studied through neuron-based and “logit lens” (nostalgebraist, 2020) methods. Neuron-based aim to identify language-specific neurons within LLMs and modify these neurons to assess their impact on the corresponding language (Zhang et al., 2024b; Zhao et al., 2024; Tang et al., 2024; Kojima et al., 2024). For example, Zhang et al. (2024b) discover removing certain neurons in LLMs leads to a significant performance decrease in some languages. Zhao et al. (2024) introduce PLND to identify activated neurons for inputs in different languages, and hypothesize that in the intermediate layers, LLMs employ English for thinking. Moreover, Tang et al. (2024); Kojima et al. (2024) explore methods to identify language-specific neurons within LLMs. However, these methods can be complex and unreliable due to “superposition,” (Elhage et al., 2022) where multiple concepts can be encoded in a single neuron. “Logit lens” methods derive token distributions from intermediate layers using the output layer’s unembedding matrix. Wendler et al. (2024) find that Llama2 (Touvron et al., 2023) might use English as an internal language in intermediate layers, and Zhong et al. (2024) extend the conclusion, showing that LLMs with continued pre-training in Japanese employ both Japanese and English in intermediate layers. Due to the varying distribution of residual streams across different layers, the “logit lens” method often has significant errors except in the last few layers, making the analysis sometimes unreliable.

### v2

@@ Abstract
The mechanisms behind multilingual capabilities in Large Language Models (LLMs) have been examined using neuron-based or internal-activation-based methods. However, these methods often face challenges such as superposition and layer-wise activation variance, which limit their reliability. Sparse Autoencoders (SAEs) offer a more nuanced analysis by decomposing the activations of LLMs into a sparse linear combination of SAE features. We introduce a novel metric to assess the monolinguality of features obtained from SAEs, discovering that some features are strongly related to specific languages. Additionally, we show that ablating these SAE features only significantly reduces abilities in one language of LLMs, leaving others almost unaffected. Interestingly, we find some languages have multiple synergistic SAE features, and ablating them together yields greater improvement than ablating individually. Moreover, we leverage these SAE-derived language-specific features to enhance steering vectors, achieving control over the language generated by LLMs. The code is publicly available at https://github.com/Aatrox103/multilingual-llm-features.
@@
Boyi Deng1, Yu Wanthanks: Corresponding authors, Yidan Zhang2, Baosong Yang, Fuli Feng1 1University of Science and Technology of China, 2Sichuan University dengboyi@mail.ustc.edu.cn
@@ 1 Introduction
Large Language Models (LLMs) (OpenAI et al., 2024; Grattafiori et al., 2024; Qwen et al., 2025) exhibit impressive abilities in various domains such as text generation (OpenAI et al., 2024; Grattafiori et al., 2024; Xu et al., 2025), instruction following (Zhang et al., 2024a; Lou et al., 2024), and reasoning (Huang and Chang, 2023). Recently, considerable efforts have been made to enhance the multilingual capabilities of LLMs to meet the growing demand for their deployment in multilingual environments (Qin et al., 2024; Huang et al., 2025). For instance, Gemini 1.5 incorporates a variety of multilingual data in its training process and emphasizes its multilingual capabilities (Team et al., 2024a)Yang et al. (2024) claim that Qwen2 supports over 30 languages and achieves great performance on multilingual benchmarks. Moreover, multilingual training data comprises approximately 3% of the training data for Llama 3, and there are also high-quality multilingual instruction-tuning data for 8 languages (Grattafiori et al., 2024). As the significance of multilingual capabilities in LLMs continues to grow, it is crucial to delve into the mechanisms of these capabilities to enhance them further.
@@
Given the advantages of SAEs, we use them to analyze multilingual capabilities in LLMs. Concretely, we start with a preliminary experiment in which we find high activation of some features in a certain language. Inspired by this, we propose a metric to measure the monolinguality of a feature based on the activation difference across different languages. The results show that some features possess strong monolingual characteristics. Moreover, we believe that these language-specific features are not only related to language-specific tokens, so we experiment on a “code-switching” (Kuwanto et al., 2024; Winata et al., 2023) dataset and find that language-specific features are also closely associated with the language-specific linguistic context. Furthermore, we use directional ablation (Arditi et al., 2024; Ferrando et al., 2024) to “zero out” language-specific features during the forward pass of LLMs, resulting in a loss of capabilities in only certain language. Interestingly, we observe that some languages may exhibit more than one specific feature. And these features have a synergistic relationship, meaning ablating these features together results in a significant improvement compared to ablating them individually.
The language-specific features we find are of great monolinguality, so we further leverage them to improve steering vectors (Turner et al., 2024). Concretely, we use language-specific features as gating signals to control steering vectors and achieve better control over the language generated by LLMs, which validates the practical potential of these language-specific features.
@@
  • We find that ablating language-specific features only significantly decreases the language-specific capabilities of LLMs.
@@ 2 Preliminary
Figure 1: The values of $\nu$, as referenced in Eq. 4, where a larger $\nu$ indicates stronger monolingualism, are reported for the top-4 features and a random feature across various languages in layer 20 of Gemma 2 2B. The values of $\nu$ for the top-4 features are greater than those of a random feature. In most languages, the top-1 feature possesses a significantly larger $\nu$. Additional results for other layers and LLMs are in Appendix C, exhibiting similar patterns. The value of the random feature (feature 2000) is too small to be visible.
@@
Flores-200 (Costa-jussà et al., 2022; Goyal et al., 2022) is a parallel corpus that contains translations of English sentences into 200 different languages. Due to the semantic similarity of the translated sentences, this dataset is particularly useful for comparing linguistic features across languages. We extract a subset called Flores-10, which includes 10 languages [^2] English (en), Spanish (es), French (fr), Japanese (ja), Korean (ko), Portuguese (pt), Thai (th), Vietnamese (vi), Chinese (zh), and Arabic (ar)..
Figure 2: The mean activation of feature 13788 across different languages in layer 10 of Gemma 2 2B. The high mean activation in Chinese suggests that feature 13788 might be related to Chinese.
@@ 3.1 Finding Language-Specific Features
To find language-specific features, we conduct a preliminary experiment by prompting Flores-10 into the LLMs and analyzing the residual stream using SAEs. We find that the mean activation of some features is particularly high for a certain language, while remaining very low for other languages, as illustrated by the example in Figure 2. Inspired by this, we propose a metric to measure the monolinguality of a feature. Specifically, given a set $\mathcal{D}=\{\mathcal{D}_{1},\ldots,\mathcal{D}_{K}\}$, which contains the residual stream set for a certain layer with $K$ different languages, we calculate the mean activation difference of feature $s$ for a specific language $L$ compared to the other languages as follows:
@@
where $\mathbf{f}_{s}(\mathbf{x})$ is the activation of feature $s$. We calculate $\nu$ for all languages and features and rank them from high to low for each language. The top-ranked features are considered language-specific features.
@@ 3.2 Monolinguality Analysis
We use the first 100 data points in Flores-10 to calculate $\nu$ for each language. The results are shown in Figure 1. From this figure, we make the following observations. (1) The mean activation of the top-4 features is significantly higher than that of a random feature, which remains close to zero. (2) For most languages, the mean activation of the top features decreases rapidly among the first few, and the mean activation of the rank #1 feature is considerably higher than the others. (3) In some languages, the rank #2 feature also shows a substantially large mean activation compared to other features. These results suggest that top-ranked features possess strong monolingual characteristics, and in most scenarios, the top-1 feature suffices in capturing these characteristics. [^3] English is the primary language for most LLMs, and it often exhibits different characteristics compared to other languages (Qin et al., 2024), so we only focus on non-English results in the subsequent sections.
@@
Figure 4: The mean activation values for the French and Korean features with various noun and prefix combinations. Introducing a different language prefix decreases the original language feature activation of nouns.
Figure 5: The changes in CE loss on texts in the target language and texts in other languages after ablating language-specific features. Ablating language-specific features has a much larger impact on the CE loss of texts in the target language compared to texts in other languages. We provide results for Gemma 2 2B here, additional results can be found in Appendix D.
@@ 4 Language-Specific Features Extend Beyond Language-Specific Tokens
In earlier sections, we only evaluate language-specific features on monolingual texts. This raises a natural question: are these language-specific features solely related to language-specific tokens? To explore this, we focus on a phenomenon called “code-switching.” [^4] Code-switching refers to the practice of alternating between two or more languages within a single text (Kuwanto et al., 2024; Winata et al., 2023). Our findings indicate that language-specific features are also related to language-specific linguistic context.
@@ 4.2 Results
We analyze the mean activation values of the Spanish features for Spanish, French, and Korean nouns, comparing scenarios with and without Spanish prefixes, as illustrated in Figure 3. Our observations are as follows: (1) Introduction of a Spanish prefix to a French or Korean noun results in higher Spanish feature activation values compared to when the French or Korean nouns stand alone. However, the value is still lower than that of the combination of Spanish prefixes and Spanish nouns. (2) The activation value for Spanish features of stand-alone French and Korean nouns remains relatively low across all layers. (3) Both French and Korean nouns with a Spanish prefix show greater increases in Spanish feature activations at deeper layers than at shallower ones. (4) Adding a Spanish prefix results in a larger increase in the Spanish feature for French nouns compared to Korean nouns. These findings suggest that adding a Spanish prefix enhances the Spanish feature activation for non-Spanish nouns, enabling the LLM to process them as if they were “Spanish tokens.” Consequently, this allows the LLM to use these non-Spanish tokens within a consistent language context.
@@
We also analyze the mean activation values of the French and Korean features for corresponding nouns, comparing scenarios with and without Spanish prefixes, as presented in the provided Figure 4. Our observations are as follows: (1) For French and Korean nouns, the original language feature activation is significantly higher when the nouns are standalone than when preceded by a Spanish prefix. (2) Both French and Korean nouns show greater decreases in their original language feature activations at deeper layers than at shallower ones. (3) Adding a Spanish prefix results in a larger decrease in the corresponding feature for French nouns compared to Korean nouns. These findings reveal that introducing a different language prefix decreases the original language feature activation of nouns, making them less like nouns from their original language.
@@
The results in Figures 3 and 4 suggest that language-specific features are not solely tied to specific language tokens but are also closely associated with the language-specific linguistic context. This suggests that the linguistic characteristics recognized by the model extend beyond individual words to encompass the contextual environment in which these words appear. Notably, the influence of a Spanish prefix is more pronounced on French nouns than on Korean nouns, potentially due to the linguistic similarities between Spanish and French. This highlights the model’s ability to dynamically adapt its feature activations based on the surrounding linguistic context, effectively reinterpreting non-Spanish tokens within a Spanish framework while diminishing their original language attributes.
@@ 5 Ablating Language-Specific Features Leads to Language-Specific Changes
Figure 6: The change in CE loss for three languages after ablating French Features. Simultaneously ablating multiple French features exhibits a synergistic effect in French, while showing no synergistic effect on other languages. We provide results for Gemma 2 2B here, and additional results can be found in Appendix F.
@@ 5.2 Ablation of Language-Specific Features
For each target language, layer, and LLM, we intervene in the inference process of LLMs using Eq. 4 on the top-2 language-specific features of the target language. We then measure the changes in CE loss for both texts in the target language and texts in other languages after ablating language-specific features. The results are shown in Figure 5. We observe that: (1) Ablating language-specific features has a much larger impact on the CE loss of texts in the target language compared to texts in other languages. (2) For different layers, the changes in CE loss of target language texts vary significantly. These findings suggest that language-specific features play a crucial role in controlling the generation process for the target language. Ablating these features from the generation process of LLMs can lead to a loss of only specific language capabilities.
@@ 5.3 Synergistic Language Features
We compare the CE loss for French, Spanish, and Japanese when using different numbers of French features for directional ablation. The results are shown in Figure 6. From this, we make the following observations: (1) In some layers, simultaneously ablating the top 2 French features for French significantly impacts the CE loss more than ablating these features individually. (2) In all layers, simultaneously ablating the top 2 French features for Spanish and Japanese results in a CE loss impact approximately equal to the sum of the effects when these features are ablated individually. (3) The changes in CE loss for French are larger than those for Spanish and Japanese. The changes for Spanish are large in some layers, while for Japanese, they are nearly zero across all layers.
@@ 6 Enhancing Steering Vectors Using Language-Specific Features
Having studied the basic characteristics of language-specific features, we now explore how to leverage these features in practice. Concretely, we use language-specific features as signals to guide steering vector (Turner et al., 2024; Rimsky et al., 2024; Mayne et al., 2024), in order to control the language in the model.
@@ 6.2 Methods
Figure 8: The y-axis shows the CE loss of LLMs identifying text in the original language as if it were in the target language, with lower values indicating better performance. The x-axis shows the impact on non-original language texts, with lower values indicating less impact. “SV” is the “steering vector” method, while “SAE” is our enhanced method. The suffix $k$L indicates results from modifying $k$ consecutive layers. Our method (green) provides a better balance between the metrics.
@@
Table 1: The results of Cross-Lingual Continuation task. Our SAE method can surpass the SV method across both metrics in most cases and achieve a much better balance between the metrics.
@@ 6.3 Results
In Figure 8, we present the results for Adversarial Language Identification. The experiments cover the steering vector both with and without SAE across one, two, and three consecutive layers, denoted as ${1L}$, ${2L}$, and ${3L}$. From the figure, we make the following observations: (1) As the number of modified layers increases, our SAE method achieves better performance in Adversarial Language Identification, whereas the performance of the SV method decreases. This indicates that SAE is more effective at Adversarial Language Identification and more robust when applied to multiple layers. (2) As the number of modified layers increases, the CE loss on other languages also increases. However, the rate of increase with our SAE method is much smaller than with the SV method. In most cases, the CE loss on other languages for SAE applied to three consecutive layers is even lower than that for SV applied to a single layer. These results suggest that our SAE method achieves a better balance between the two metrics on Adversarial Language Identification.
@@ 7 Related Works
Multilingual mechanisms of LLMs are mainly studied through neuron-based and “logit lens” (nostalgebraist, 2020) methods. Neuron-based aim to identify language-specific neurons within LLMs and modify these neurons to assess their impact on the corresponding language (Zhang et al., 2024b; Zhao et al., 2024; Tang et al., 2024; Kojima et al., 2024). For example, Zhang et al. (2024b) discover that removing certain neurons in LLMs leads to a significant performance decrease in some languages. Zhao et al. (2024) introduce PLND to identify activated neurons for inputs in different languages, and hypothesize that in the intermediate layers, LLMs employ English for thinking. Moreover, Tang et al. (2024); Kojima et al. (2024) explore methods to identify language-specific neurons within LLMs. However, these methods can be complex and unreliable due to “superposition,” (Elhage et al., 2022) where multiple concepts can be encoded in a single neuron. “Logit lens” methods derive token distributions from intermediate layers using the output layer’s unembedding matrix. Wendler et al. (2024) find that Llama2 (Touvron et al., 2023) might use English as an internal language in intermediate layers, and Zhong et al. (2024) extend the conclusion, showing that LLMs with continued pre-training in Japanese employ both Japanese and English in intermediate layers. Due to the varying distribution of residual streams across different layers, the “logit lens” method often has significant errors except in the last few layers, making the analysis sometimes unreliable.